# Two-component relativistic equation-of-motion coupled cluster for electron ionization

### v1

@@ I Introduction
An alternative strategy that still falls within the CC/EOMCC family is to begin with a reference state that is well-described by a SR method but differs from the target open-shell state by particle number. For example, the ionization potential (IP) EOMCC approach[34, 35, 36, 37, 38, 39, 40, 41, 42] begins with a closed-shell $N$-electron state, which should be well-described by the CC hierarchy applied to a restricted Hartree–Fock (RHF) reference configuration, and parametrizes the open-shell $(N-1)$-electron state (the ionized state) using non-particle-conserving excitation operators that remove an electron from the reference state. The key benefit of this approach is that, by starting from properly spin-adapted reference wave function that does not break desired orbital degeneracies, the subsequent open-shell energetics should retain this property. This advantage becomes more apparent when considering energy splittings due to spin–orbit coupling (SOC) effects in relativistic calculations, where spin-contamination-induced degeneracy issues could lead to energy errors on the order of the splittings themselves. Moreover, even though $S_{z}$ and $S^{2}$ are technically not good quantum numbers upon the introduction of SOC, one can still take advantage of time-reversal symmetry in the form of Kramers pairs for closed-shell determinants, which in practice produces degenerate pairs of spinors as one would obtain in the non-relativistic RHF case.[43, 44, 45] This symmetry preservation extends the benefits of IP-EOMCC to the relativistic domain, as the underlying structure of relativistic CC calculations would be analogous to that of non-relativistic singlet spin-adapted CC.[46]
@@
This work uses CC/IP-EOMCC to examine ionization potentials of HCCX (X = Cl, Br, I) systems, which have been characterized both experimentally and theoretically, from as far back as the 1970s.[47, 48, 49, 50, 51, 52, 53, 54, 55, 56] More recently, the theoretical study presented in Ref. 57 provided great details about the geometries, vibrational frequencies, and both vertical and adiabatic IP values of these molecules. However, that study did not account for relativistic effects, which are necessary to describe SOC-induced splitting of the doublet ground and excited states of HCCX+. Here, we apply relativistic CC/IP-EOMCC to this problem, and the relativistic treatment we employ falls within the exact two-component (X2C) [58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81] family of methods. There are a number of prior studies that extend the IP-EOMCC approach to the relativistic domain using X2C[82, 83, 84, 85, 86] or other[87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97] relativistic frameworks. Focusing on the X2C-based calculations, the majority of these investigations used spin-free X2C, and the most sophisticated calculations have considered only the Dirac–Coulomb (DC) or Dirac–Coloumb–Gaunt (DCG) Hamiltonians, as opposed to the full Dirac–Coulomb–Breit (DCB) Hamiltonian.[84, 86] As for correlation effects, to the best of our knowledge, only one of these studies has gone beyond the IP-EOMCCSD level of theory; Ref. 85 treated K-edge ionization features at up to the IP-EOMCCSDTQ level, but it should be noted that these high-order correlation effects were assessed only within a spin-free X2C formalism. The present study aims to assess high-order correlation effects (up to triple excitations in both the ground-state CC and IP-EOM parts of the algorithm), as well as a mean-field treatment of the full DCB Hamiltonian through the molecular mean-field (mmf)[98, 99, 100, 79] X2C formalism.
@@ II.1 The X2C relativistic framework
The relativistic frameworks used in this work are the mmf formulation of X2C and a more approximate one-electron X2C approach (1eX2C). Detailed comparisons of mmfX2C and 1eX2C can be found in Refs. 79 and 98 and the references contained therein. In this Section, we only summarize the details that are relevant to this work.
@@
In terms of the implementation, we recall that the 4c expansion of any atomic-orbital integrals should be done in a fully uncontracted basis to prevent variational collapse or prolapse of the 4c DHF solution.[105, 106, 107, 108, 109, 102, 103, 104] However, after performing the X2C transformation, one can recontract the 2c integrals using the basis set’s original contraction coefficients.[79] Given that we are working in a 2c space only, the post-SCF calculations also employ the no-virtual-pair approximation,[110, 111, 112, 113, 114] in which electron–positron correlation effects are ignored.
@@ II.2 IP-EOMCC theory
By setting $M^{\prime}=M$ in Eq. (13), one arrives at the usual IP-EOMCCSD ($M^{\prime}=M=2$), IP-EOMCCSDT ($M^{\prime}=M=3$), etc. methods, but, again, $M^{\prime}$ and $M$ need not be equal in general. As an example, in the IP-EOMCCSD(3h2p)[41, 42] approach, one diagonalizes the CCSD ($M=2$) similarity-transformed Hamiltonian in the space spanned by |$\ket*{\Phi_{i}}$, |$\ket*{\Phi_{ij}^{\phantom{i}b}}$, and |$\ket*{\Phi_{ijk}^{\phantom{i}bc}}$ determinants ($M^{\prime}=3$), which refer to configurations that are generated from $|\Phi_{0}\rangle$ by the removal of an electron, plus up to two-particle transitions. It has been shown[41, 42] that the correlation effects folded into the CCSD $\bar{H}$, i.e., excluding $\hat{T}_{3}$, may be sufficient for IP-EOMCCSD(3h2p) to describe the multi-reference character of open-shell radicals, even though the 3h2p transitions correspond to parts of triple excitations in the ($N-1$)-electron manifold. This mismatch in excitation level may overcorrelate the ionized states relative to the $N$-electron ground state, leading to absolute IP values that are too low. However, the electronic spectrum of the $(N-1)$-electron species relative to the $K=1$ state may be improved compared to that determined via IP-EOMCCSD.[41, 42] IP-EOMCCSD(3h2p) is also less computationally demanding than full IP-EOMCCSDT due to the neglect of $\hat{T}_{3}$ in the ground-state CC part of the algorithm and in $\bar{H}$. Additional details about the IP-EOMCCSD(3h2p) approach can be found in Refs. 41, 42.
@@ III Computational Details
Section IV begins with a study of the effect that the basis set has on vertical IPs to the X and A ${}^{2}\Pi_{\Omega}$ ($\Omega={\nicefrac{{3}}{{2}}}$ and ${\nicefrac{{1}}{{2}}}$) states computed at the IP-EOMCCSD level of theory. For this study, structures for neutral HCCCl and HCCBr were taken from NIST’s Computational Chemistry Comparison and Benchmark Database,[115] whereas the structure for HCCI was taken from experimental data reported in Ref. 53. The structural parameters are reproduced in Table 1. The basis sets examined were the ANO-RCC-VDZP, ANO-RCC-VTZP, and ANO-RCC-VQZP basis sets,[116] the x2c-SVPall-2c, x2c-TZVPPall-2c, and x2c-QZVPPall-2c sets,[117, 118] and the Dyall v2z, v3z, and v4z sets.[119, 120, 121, 122, 123] The ANO-RCC- and x2c-type basis sets were taken from the Basis Set Exchange,[124, 125, 126] while the uncontracted Dyall bases were obtained from Ref. 123.
@@
The accuracy of X2C-based IP-EOMCC was also evaluated against the adiabatic data for the haloacetylene molecules reported in Ref. 48. For this study, the geometries for each of the molecules are taken from Ref. 57, which were optimized using the cc-pVTZ[127, 128, 129, 130] basis set at the non-relativistic CCSD level of theory, in the case of the neutral species, and non-relativistic IP-EOMCCSD, in the case of the cation states. The resulting geometric parameters are reproduced in Table 2. In the adiabatic IP calculations, where we focus on the DCB-X2C relativistic framework, we employed IP-EOMCCSD using the ANO-RCC-VDZP and ANO-RCC-VQZP basis sets, as well as IP-EOMCCSD(3h2p) and IP-EOMCCSDT calculations in the ANO-RCC-VDZP basis.
Table 2: Equilibrium bond lengths of the HCCX molecules in their ground state and the first- and second-ionized states reported in Ref. 57.
@@ IV.1.1 HCCCl
Expt (vertical) [^1] Vertical transition from Ref. 47. The experimental IP values for the X and A ${}^{2}\Pi_{\nicefrac{{1}}{{2}}}$ states are not available. 10.63 14.08
Expt (adiabatic) [^2] Adiabatic transition from Ref. 48. 10.58 10.60 0.02 13.87 13.92 0.05
As shown in Fig. 1 and Table 3, IP-EOMCCSD underestimates the experimentally determined vertical IP data by about 0.1 eV for both the X and A ${}^{2}\Pi_{\nicefrac{{3}}{{2}}}$ states, when combined with the x2c-SVPall-2c, ANO-RCC-VDZP, or Dyall v2z basis sets. The calculated vertical IPs increase by about 0.2 eV when the x2c-TZVPPall-2c, ANO-RCC-VTZP, and Dyall v3z sets are employed and by another 0.1 eV when using the x2c-QZVPPall-2c, ANO-RCC-VQZP, and Dyall v4z basis sets. For the X ${}^{2}\Pi_{\nicefrac{{3}}{{2}}}$ state of HCCCl+, the double-$\zeta$ basis sets provide the best estimates of the experimentally obtained value to within 0.7 eV or less, with 1eX2C-IP-EOMCCSD/ANO-RCC-VDZP and DCB-X2C-IP-EOMCCSD/x2c-SVPall-2c being the outliers having 0.10–0.014 eV deviations relative to experimental data. In the case of the A ${}^{2}\Pi_{\nicefrac{{3}}{{2}}}$ state, the triple-$\zeta$ basis sets provide the best estimate of the experimentally determined vertical IP value with errors in the range of 0.03–0.06 eV. It is also interesting to note that the calculated $\Delta_{\Omega}$ splittings characterizing both the X and A ion states are essentially independent of the basis set and the relativistic treatment (see Table 3). Both 1eX2C- and DCB-X2C-IP-EOMCCSD predict the experimentally measured splitting values to within 0.01 eV, which is consistent with the experimental uncertainty range.[48]
@@ IV.1.2 HCCBr
Computed and experimentally-obtained vertical IPs for HCCBr are provided in Fig. 2 and Table 4. As can be seen in Table 4, more complete vertical IP data are provided in this case, as compared to HCCCl. The data in Table 4 also demonstrate that experimentally-obtained vertical and adiabatic IP values are in reasonable agreement, to within 0.06–0.07 eV. However, we suspect that the adiabatic IP values are more reliable than the vertical IP values, for the following reasons. First, the experimental data for the X ${}^{2}\Pi_{\Omega}$ states seem to be inconsistent because the adiabatic IP values are larger than the vertical ones, whereas geometric relaxation effects should lead to the opposite trend. Second, Ref. 48 notes challenges in assigning the vertical IPs for the A ${}^{2}\Pi_{\nicefrac{{3}}{{2}}}$ and ${}^{2}\Pi_{\nicefrac{{1}}{{2}}}$ states because the splitting is similar in magnitude to the vibrational spacings for the molecules.
@@
As for the computed data, we find once again that the vertical IP values are underestimated when using small basis sets, and the predicted IP values increase as the basis set improves from double- to triple-$\zeta$ quality (0.20–0.25 eV for the X ${}^{2}\Pi_{\Omega}$ states) and from triple- to quadruple-$\zeta$ quality (0.10–0.15 eV, again, for the X ${}^{2}\Pi_{\Omega}$ states). These increases are slightly larger than those we observed for HCCCl. The calculated vertical IP values for the A ${}^{2}\Pi_{\Omega}$ states produce slightly smaller errors (cf. Table 4). Similar to the HCCCl molecule, the double-$\zeta$ basis sets produce the best vertical IP estimates relative to experimentally determined values for both X and A ${}^{2}\Pi_{\Omega}$ states. As was the case for HCCCl, the good performance of small basis sets is simply due to a cancellation of errors, and the large errors associated with calculations carried out in quadruple-$\zeta$ basis sets is mainly due to correlation effects that are missing at the IP-EOMCCSD level of theory.
@@ IV.1.3 HCCI
Computed and experimentally-obtained vertical IPs for HCCI are provided in Fig. 3 and Table 5. Of the three cations studied in this work, HCCI+ is characterized by the largest SOC splittings in the X ${}^{2}\Pi_{\Omega}$ ($\sim$0.4 eV) and A ${}^{2}\Pi_{\Omega}$ ($\sim$0.25 eV) states. As such, this system serves as a good benchmark for approaches that account for spin-dependent relativistic effects. Like HCCBr, the experimental data for the vertical and adiabatic IPs of HCCI are more complete than for HCCCl. However, unlike HCCBr, the relative magnitudes of the vertical and adiabatic IPs make physical sense for HCCI, with the adiabatic IPs having the lower values, for both the X and A ${}^{2}\Pi_{\Omega}$ states. As a result, we should be able to directly compare the quality of our calculated vertical IP data with the experimentally determined values from Ref. 47.
@@ IV.1.4 Basis set recontraction with 1eX2C and DCB-X2C
Cation State IP (recontracted) [^1] Vertical transition from Ref. 47.Vertical transition from Ref. 47. IP (uncontracted)
@@ IV.2 Correlation effects beyond IP-EOMCCSD
In the following analysis, we apply both IP-EOMCCSD and higher-order correlation treatments to the adiabatic, rather than vertical, IPs for the HCCX molecules. We focus on the adiabatic IPs because the available experimental data in Ref. 48 are more recent and complete than those for the vertical IPs reported in Ref. 47. We also limit our study to DCB-X2C-based treatments of relativistic effects in order to minimize possible sources of error. In the interest of computational efficiency, we wish to evaluate the $\mathscr{O}(\mathscr{N}^{8})$-scaling CCSDT and $\mathscr{O}(\mathscr{N}^{7})$-scaling IP-EOMCCSDT/IP-EOMCCSD(3h2p) portions of the algorithm using a contracted basis set. From our vertical IP analysis, we recall that, of the two contracted basis set families, the ANO-RCC family provides the best agreement with results obtained using Dyall-type basis sets when the relativistic treatment is DCB-X2C. As such, we do not consider the x2c-type basis sets below.
@@
The equilibrium geometries of the neutral and two lowest ion states, as summarized in Table 2, all have $C_{\infty v}$ symmetry.[57] The vibrational analysis in Ref. 57 indicates that the difference between the zero-point vibrational energies of the neutral molecule and cation in either the X ${}^{2}\Pi_{\Omega}$ and A ${}^{2}\Pi_{\Omega}$ states are on the order of only 100 cm-1 or around 0.01 eV. Given that the experimental uncertainty for the adiabatic 0–0 transition reported in Ref. 48 is also 0.01 eV, the computed electronic A–X adiabatic transition energy, $\omega_{\mathrm{A-X}}$, should be an acceptable approximation to the experimentally-obtained value.
Figure 4: The calculated adiabatic IPs for HCCCl, HCCBr, and HCCI using DCB-X2C-IP-EOMCCSD/ANO-RCC-VQZP (blue) and the composite schemes described in Eqs. (16) (orange) and (17) (green). The experimentally determined IPs from Ref. 48 are shown as black dotted lines. The states ordering, from the lowest to highest energy, for each molecule is X ${}^{2}\Pi_{\nicefrac{{3}}{{2}}}$, X ${}^{2}\Pi_{\nicefrac{{1}}{{2}}}$, A ${}^{2}\Pi_{\nicefrac{{3}}{{2}}}$, and A ${}^{2}\Pi_{\nicefrac{{1}}{{2}}}$.
@@
Experiment [^2] Adiabatic transition from Ref. 48.Adiabatic transition from Ref. 48. 10.58 10.60 0.02 13.87 13.92 0.05 26540 26780

### v2

@@ I Introduction
An alternative strategy that still falls within the CC/EOMCC family is to begin with a reference state that is well-described by a SR method but differs from the target open-shell state by particle number. For example, the ionization potential (IP) EOMCC approach[34, 35, 36, 37, 38, 39, 40, 41, 42, 43] begins with a closed-shell $N$-electron state, which should be well-described by the CC hierarchy applied to a restricted Hartree–Fock (RHF) reference configuration, and parametrizes the open-shell $(N-1)$-electron state (the ionized state) using non-particle-conserving excitation operators that remove an electron from the reference state. The key benefit of this approach is that, by starting from properly spin-adapted reference wave function that does not break desired orbital degeneracies, the subsequent open-shell energetics should retain this property. This advantage becomes more apparent when considering energy splittings due to spin–orbit coupling (SOC) effects in relativistic calculations, where spin-contamination-induced degeneracy issues could lead to energy errors on the order of the splittings themselves. Moreover, even though $S_{z}$ and $S^{2}$ are technically not good quantum numbers upon the introduction of SOC, one can still take advantage of time-reversal symmetry in the form of Kramers pairs for closed-shell determinants, which in practice produces degenerate pairs of spinors as one would obtain in the non-relativistic RHF case.[44, 45, 46] This symmetry preservation extends the benefits of IP-EOMCC to the relativistic domain, as the underlying structure of relativistic CC calculations would be analogous to that of non-relativistic singlet spin-adapted CC.[47]
@@
This work uses CC/IP-EOMCC to examine ionization potentials of HCCX (X = Cl, Br, I) systems, which have been characterized both experimentally and theoretically, from as far back as the 1970s.[48, 49, 50, 51, 52, 53, 54, 55, 56, 57] More recently, the theoretical study presented in Ref. 58 provided great details about the geometries, vibrational frequencies, and both vertical and adiabatic IP values of these molecules. However, that study did not account for relativistic effects, which are necessary to describe SOC-induced splitting of the doublet ground and excited states of HCCX+. Here, we apply relativistic CC/IP-EOMCC to this problem, and the relativistic treatment we employ falls within the exact two-component (X2C) [59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82] family of methods. There are a number of prior studies that extend the IP-EOMCC approach to the relativistic domain using X2C[83, 84, 85, 86, 87] or other[88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98] relativistic frameworks. Focusing on the X2C-based calculations, the majority of these investigations used spin-free X2C, and the most sophisticated calculations have considered only the Dirac–Coulomb (DC) or Dirac–Coloumb–Gaunt (DCG) Hamiltonians, as opposed to the full Dirac–Coulomb–Breit (DCB) Hamiltonian.[85, 87] As for correlation effects, to the best of our knowledge, only one of these studies has gone beyond the IP-EOMCCSD level of theory; Ref. 86 treated K-edge ionization features at up to the IP-EOMCCSDTQ level, but it should be noted that these high-order correlation effects were assessed only within a spin-free X2C formalism. The present study aims to assess high-order correlation effects (up to triple excitations in both the ground-state CC and IP-EOM parts of the algorithm), as well as a mean-field treatment of the full DCB Hamiltonian through the molecular mean-field (mmf)[99, 100, 85, 80] X2C formalism.
@@ II.1 The X2C relativistic framework
The relativistic frameworks used in this work are the mmf formulation of X2C and a more approximate one-electron X2C approach (1eX2C). Detailed comparisons of mmfX2C and 1eX2C can be found in Refs. 80 and 99 and the references contained therein. In this Section, we only summarize the details that are relevant to this work.
@@
In terms of the implementation, we recall that the 4c expansion of any atomic-orbital integrals should be done in a fully uncontracted basis to prevent variational collapse or prolapse of the 4c DHF solution.[105, 106, 107, 108, 109, 102, 103, 104] However, after performing the X2C transformation, one can recontract the 2c integrals using the basis set’s original contraction coefficients.[80] Given that we are working in a 2c space only, the post self-consistent field (SCF) calculations also employ the no-virtual-pair approximation,[110, 111, 112, 113, 114] in which electron–positron correlation effects are ignored.
@@ II.2 IP-EOMCC theory
By setting $M^{\prime}=M$ in Eq. (13), one arrives at the usual IP-EOMCCSD ($M^{\prime}=M=2$), IP-EOMCCSDT ($M^{\prime}=M=3$), etc. methods, but, again, $M^{\prime}$ and $M$ need not be equal in general. As an example, in the IP-EOMCCSD(3h2p)[41, 42] approach, one diagonalizes the CCSD ($M=2$) similarity-transformed Hamiltonian in the space spanned by |$\ket*{\Phi_{i}}$, |$\ket*{\Phi_{ij}^{\phantom{i}b}}$, and |$\ket*{\Phi_{ijk}^{\phantom{i}bc}}$ determinants ($M^{\prime}=3$), which refer to configurations that are generated from $|\Phi_{0}\rangle$ by the removal of an electron, plus up to two-particle transitions. It has been shown[41, 42] that the correlation effects folded into the CCSD $\bar{H}$, i.e., excluding $\hat{T}_{3}$, may be sufficient for IP-EOMCCSD(3h2p) to describe the multi-reference character of open-shell radicals, even though the 3h2p transitions correspond to parts of triple excitations in the ($N-1$)-electron manifold. This mismatch in excitation level may overcorrelate the ionized states relative to the $N$-electron ground state, leading to absolute IP values that are too low. However, the electronic spectrum of the $(N-1)$-electron species relative to the $K=1$ state may be improved compared to that determined via IP-EOMCCSD.[41, 42] IP-EOMCCSD(3h2p) is also less computationally demanding than full IP-EOMCCSDT due to the neglect of $\hat{T}_{3}$ in the ground-state CC part of the algorithm and in $\bar{H}$. Additional details about the IP-EOMCCSD(3h2p) approach can be found in Refs. 41, 42, 43.
@@ III Computational Details
[…]
@@
The accuracy of X2C-based IP-EOMCC was also evaluated against the adiabatic data for the haloacetylene molecules reported in Ref. 49. For this study, the geometries for each of the molecules are taken from Ref. 58, which were optimized using the cc-pVTZ[128, 129, 130, 131] basis set at the non-relativistic CCSD level of theory, in the case of the neutral species, and non-relativistic IP-EOMCCSD, in the case of the cation states. The resulting geometric parameters are reproduced in Table 2. In the adiabatic IP calculations, where we focus on the DCB-X2C relativistic framework, we employed IP-EOMCCSD using the ANO-RCC-VDZP and ANO-RCC-VQZP basis sets, as well as IP-EOMCCSD(3h2p) and IP-EOMCCSDT calculations in the ANO-RCC-VDZP basis.
Table 2: Equilibrium bond lengths of the HCCX molecules in their ground state and the first- and second-ionized states reported in Ref. 58.
@@ IV.1.1 HCCCl
Expt (vertical) [^1] Vertical transition from Ref. 48. The experimental IP values for the X and A ${}^{2}\Pi_{\nicefrac{{1}}{{2}}}$ states are not available. 10.63 14.08
Expt (adiabatic) [^2] Adiabatic transition from Ref. 49. 10.58 10.60 0.02 13.87 13.92 0.05
As shown in Fig. 1 and Table 3, IP-EOMCCSD underestimates the experimentally determined vertical IP data by about 0.1 eV for both the X and A ${}^{2}\Pi_{\nicefrac{{3}}{{2}}}$ states, when combined with the x2c-SVPall-2c, ANO-RCC-VDZP, or Dyall v2z basis sets. The calculated vertical IPs increase by about 0.2 eV when the x2c-TZVPPall-2c, ANO-RCC-VTZP, and Dyall v3z sets are employed and by another 0.1 eV when using the x2c-QZVPPall-2c, ANO-RCC-VQZP, and Dyall v4z basis sets. For the X ${}^{2}\Pi_{\nicefrac{{3}}{{2}}}$ state of HCCCl+, the double-$\zeta$ basis sets provide the best estimates of the experimentally obtained value to within 0.7 eV or less, with 1eX2C-IP-EOMCCSD/ANO-RCC-VDZP and DCB-X2C-IP-EOMCCSD/x2c-SVPall-2c being the outliers having 0.10–0.014 eV deviations relative to experimental data. In the case of the A ${}^{2}\Pi_{\nicefrac{{3}}{{2}}}$ state, the triple-$\zeta$ basis sets provide the best estimate of the experimentally determined vertical IP value with errors in the range of 0.03–0.06 eV. It is also interesting to note that the calculated $\Delta_{\Omega}$ splittings characterizing both the X and A ion states are essentially independent of the basis set and the relativistic treatment (see Table 3). Both 1eX2C- and DCB-X2C-IP-EOMCCSD predict the experimentally measured splitting values to within 0.01 eV, which is consistent with the experimental uncertainty range.[49] It is also worth noting that, as shown in Table S2 in the Supplemental Information, the IP values obtained using the cc-pVDZ-DK through cc-pVQZ-DK basis sets are very close to their analogs reported in Table 3. Focusing on the cc-pVQZ-DK results, the IP-EOMCCSD energetics for HCCCl are within 0.01–0.02 eV of the values evaluated using the ANO-RCC-VQZP, Dyall v4z, and x2c-QZVPPall-2c basis sets. Furthermore, the use of the cc-pV5Z-DK basis set increases the IP values by $\sim$0.04–0.05 eV compared to those computing using the cc-pVQZ-DK set. The $\Delta_{\Omega}$ splittings in HCCCl are also practically unaffected by the basis set size up to cc-pV5Z-DK. Thus, the IP values of HCCCl obtained using quadruple-$\zeta$-level basis sets could be considered converged with respect to basis set size to well within 0.1 eV.
@@ IV.1.2 HCCBr
Computed and experimentally-obtained vertical IPs for HCCBr are provided in Fig. 2 and Table 4. As can be seen in Table 4, more complete vertical IP data are provided in this case, as compared to HCCCl. The data in Table 4 also demonstrate that experimentally-obtained vertical and adiabatic IP values are in reasonable agreement, to within 0.06–0.07 eV. However, we suspect that the adiabatic IP values are more reliable than the vertical IP values, for the following reasons. First, the experimental data for the X ${}^{2}\Pi_{\Omega}$ states seem to be inconsistent because the adiabatic IP values are larger than the vertical ones, whereas geometric relaxation effects should lead to the opposite trend. Second, Ref. 49 notes challenges in assigning the vertical IPs for the A ${}^{2}\Pi_{\nicefrac{{3}}{{2}}}$ and ${}^{2}\Pi_{\nicefrac{{1}}{{2}}}$ states because the splitting is similar in magnitude to the vibrational spacings for the molecules.
@@
[…]
@@ IV.1.3 HCCI
Computed and experimentally-obtained vertical IPs for HCCI are provided in Fig. 3 and Table 5. Of the three cations studied in this work, HCCI+ is characterized by the largest SOC splittings in the X ${}^{2}\Pi_{\Omega}$ ($\sim$0.4 eV) and A ${}^{2}\Pi_{\Omega}$ ($\sim$0.25 eV) states. As such, this system serves as a good benchmark for approaches that account for spin-dependent relativistic effects. Like HCCBr, the experimental data for the vertical and adiabatic IPs of HCCI are more complete than for HCCCl. However, unlike HCCBr, the relative magnitudes of the vertical and adiabatic IPs make physical sense for HCCI, with the adiabatic IPs having the lower values, for both the X and A ${}^{2}\Pi_{\Omega}$ states. As a result, we should be able to directly compare the quality of our calculated vertical IP data with the experimentally determined values from Ref. 48.
@@ IV.1.4 Basis set recontraction with 1eX2C and DCB-X2C
Cation State IP (recontracted) [^1] Vertical transition from Ref. 48.Vertical transition from Ref. 48. IP (uncontracted)
@@ IV.2 Correlation effects beyond IP-EOMCCSD
In the following analysis, we apply both IP-EOMCCSD and higher-order correlation treatments to the adiabatic, rather than vertical, IPs for the HCCX molecules. We focus on the adiabatic IPs because the available experimental data in Ref. 49 are more recent and complete than those for the vertical IPs reported in Ref. 48. We also limit our study to DCB-X2C-based treatments of relativistic effects in order to minimize possible sources of error. In the interest of computational efficiency, we wish to evaluate the $\mathscr{O}(\mathscr{N}^{8})$-scaling CCSDT and $\mathscr{O}(\mathscr{N}^{7})$-scaling IP-EOMCCSDT/IP-EOMCCSD(3h2p) portions of the algorithm using a contracted basis set. From our vertical IP analysis, we recall that, of the two contracted basis set families, the ANO-RCC family provides the best agreement with results obtained using Dyall-type basis sets when the relativistic treatment is DCB-X2C. As such, we do not consider the x2c-type basis sets below.
@@
The equilibrium geometries of the neutral and two lowest ion states, as summarized in Table 2, all have $C_{\infty v}$ symmetry.[58] The vibrational analysis in Ref. 58 indicates that the difference between the zero-point vibrational energies of the neutral molecule and cation in either the X ${}^{2}\Pi_{\Omega}$ and A ${}^{2}\Pi_{\Omega}$ states are on the order of only 100 cm-1 or around 0.01 eV. Given that the experimental uncertainty for the adiabatic 0–0 transition reported in Ref. 49 is also 0.01 eV, the computed electronic A–X adiabatic transition energy, $\omega_{\mathrm{A-X}}$, should be an acceptable approximation to the experimentally-obtained value.
Figure 4: The calculated adiabatic IPs for HCCCl, HCCBr, and HCCI using DCB-X2C-IP-EOMCCSD/ANO-RCC-VQZP (blue) and the composite schemes described in Eqs. (16) (orange) and (17) (green). The experimentally determined IPs from Ref. 49 are shown as black dotted lines. The states ordering, from the lowest to highest energy, for each molecule is X ${}^{2}\Pi_{\nicefrac{{3}}{{2}}}$, X ${}^{2}\Pi_{\nicefrac{{1}}{{2}}}$, A ${}^{2}\Pi_{\nicefrac{{3}}{{2}}}$, and A ${}^{2}\Pi_{\nicefrac{{1}}{{2}}}$.
@@
Experiment [^2] Adiabatic transition from Ref. 49.Adiabatic transition from Ref. 49. 10.58 10.60 0.02 13.87 13.92 0.05 26540 26780